Research Index / Hercules / Schedule of previous weeks / 2003

SCHEDULE OF PREVIOUS WEEKS - 2003

March 2003

DATE

ACTIVITY

03/03 Petawatt design, Cleaner alignment
03/04 Petawatt design
03/05 Petawatt design, regen is fixed
03/06 Petawatt design, laser maintenance
03/07 Petawatt design, laser maintenance

 

DATE

ACTIVITY

03/10 Regen in diagnostic regime for LILAC alignment (starting 10:10 AM), petawatt design
03/11 Regen in diagnostic regime for LILAC alignment (starting 10:20 AM), petawatt design
03/12 Regen in diagnostic regime for LILAC alignment, (starting 10:30 AM), laser has been switched off at 4:30 PM due problems with chiller, petawatt design
03/13 Regen has been restored at 12 AM, LILAC alignment, petawatt design
03/14 Regen in diagnostic regime for LILAC alignment, (starting 10:30 AM), petawatt design

 

DATE

ACTIVITY

03/17 Regen in diagnostic regime for wavefront correction (started 10:00AM), petawatt pump alignment
03/18 Regen in diagnostic regime for LILAC alignment (started 10:00 AM),  petawatt pump alignment
03/19 Regen in diagnostic regime for LILAC alignment (started 10:00 AM),  petawatt pump alignment
03/20 Regen in diagnostic regime for LILAC alignment (started 10:00 AM),  petawatt pump alignment
03/21 Regen in diagnostic regime for LILAC alignment (started 10:30 AM),  petawatt pump power measurements (22 J in one channel)

 

DATE

ACTIVITY

03/24 Regen in diagnostic regime for  LILAC alignment (started 10:00AM), petawatt pump energy measurements (34J in one channel). Turned off 9:00 PM.
03/25 Regen in full power regime for  LILAC alignment (started 10:00AM), compressors alignment for  LILAC, petawatt pump energy measurements (37J in one channel). Turned off 7:00 PM.
03/26 Regen in diagnostic regime for  LILAC alignment (started 10:00AM), regen in full power regime,  pulse duration measurements (25 fs) for  LILAC, cleaner is on, fairing with 10 TW (started 6:00 PM), petawatt pump energy measurements (47J in one channel). Turned off 9:30 PM.
03/27 Regen in diagnostic regime (turned on 10:00 AM) for  LILAC alignment (started 12:30 PM), regen in full power regime for  LILAC alignment and wavefront correction (started 4:30 PM). Turned off 8:15 PM.
03/28 Regen in diagnostic regime (turned on 10:10 AM) for  wavefront correction (started 3:00 PM), regen in full power (started 5:00PM).

 

April 2003

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

03/31

Monday

Regen in diagnostic regime for far field measurements (started 10:00AM), regen in full power for LILAC (started 4:00 PM).  

alignment

 

04/01

Tuesday

Regen in diagnostic regime (started 10:00AM), regen in full power. Laser maintenance (from 5:00 PM to 10:00 PM)  

alignment

 

04/02

Wednesday

 

Regen in diagnostic regime (started 10:00AM), regen in full power, cleaner on, four path amp. on (from 5:00 PM to 9:30 PM)  

alignment and firing with 10 TW

 

04/03

Thursday

Regen in diagnostic regime (started 10:00AM), regen in full power, cleaner on, four path amp. on (from 4:00 PM to 5:00PM) Chiller problems (started 5:00 PM) Alignment for wavefront correction

alignment and firing with 10 TW

 

04/04

Friday

Regen in diagnostic regime , regen in full power for laser maintenance (started 10:00)  

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

04/07

Monday

Regen in diagnostic regime  (started 12:00AM), chiller problems

alignment with He-NE

alignment

 

04/08

Tuesday

Regen in diagnostic regime (started 10:00AM), regen in full power, cleaner on, four-pass amplifier on. Was switched - off at 5:30 PM because of problems with cryogenic chamber.

alignment with He-NE

Firing with 10 TW

 

04/09

Wednesday

 

Regen in diagnostic regime , regen in full power, cleaner on, four-pass amplifier on in 1 Hz regime. Laser maintenance .

alignment with He-NE

 

petawatt pump alignment

04/10

Thursday

Regen in diagnostic regime, (started 10:00AM) regen in full power, cleaner on, four-pass amplifier on in 1 Hz regime, glass pump on. Laser maintenance 

alignment with He-NE

Firing with 30 TW

 

04/11

Friday

Regen in diagnostic regime , regen in full power for laser maintenance.

alignment with regen

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

04/14

Monday

 

chiller problems

 

alignment with He-NE

   

04/15

Tuesday

 

chiller problems

 

alignment with He-NE

 

 

04/16

Wednesday

 

 

chiller problems, laser maintenance

 

alignment with He-NE

 

 

04/17

Thursday

 

chiller problems

 

alignment with He-NE

 

 

04/18

Friday

chiller problems

 

alignment with regen

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

04/21

Monday

 

Regen in full power, laser maintenance

 

 

 

04/22

Tuesday

 

Regen in diagnostic regime, laser maintenance

 

alignment

 

 

04/23

Wednesday

 

 

Regen in diagnostic regime, laser maintenance

 

alignment

 

 

04/24

Thursday

 

Chiller problem

 

alignment

 

 

04/25

Friday

Chiller problem

 

alignment

 

   

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

04/28

Monday

 

Regen in full power, cleaner on, 4-pass on, laser maintenance

 

 

 

04/29

Tuesday

 

Regen in full power, cleaner on, 4-pass on, laser maintenance

 

 

 

 

04/30

Wednesday

 

 

Regen in full power, cleaner on, 4-pass on, laser maintenance

 

 

 

 

04/01

Thursday

 

Regen in full powe (started 10:00 AM)r, cleaner on, 4-pass on, 2-pass on (off 9:00 PM).

 

 

Firing 30 TW

 

04/02

Friday

Regen in full power.

 

Wavefront measurement preparation, alignment

 

   

 

May 2003

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

05/04

Sunday

 

Regen in full power, cleaner on, four-pass on, two-pass on, laser power measurements

 

 

 

05/05

Monday

 

Regen in diagnostic regime, laser maintenance -

wavefront correction alignment

 

 

05/06

Tuesday

 

 

Regen in full power, cleaner on, four-pass on, two-pass on.

Firing with 30 TW

 

 

05/07

Wednesday

 

Regen in diagnostic regime , laser maintenance

 

 

 

 

05/08

Thursday

Regen in full power, cleaner on, four-pass on, two-pass on

 

 

 

alignment

 

 

 

05/09

Friday

Regen in diagnostic regime , laser maintenance: four-pass pump relay recover

 

 

 

 

   

 

 

05/10

Saturday

Regen in full power, cleaner on, four-pass on, two-pass on

 

 

 

alignment

 

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

05/13

Monday

 

Regen in diagnostic regime, laser maintenance

 

 

 

05/14

Tuesday

 

 

Regen in diagnostic regime , laser maintenance:the test of damage threshold for diffraction grating samples from Lawrence Livermore lab.

 

 

 

05/15

Wednesday

 

Regen in diagnostic regime , laser maintenance:the test of damage threshold for diffraction grating samples from Lawrence Livermore lab.

 

 

pump rod alignment

05/16

Thursday

Regen in diagnostic regime , laser maintenance:the test of damage threshold for diffraction grating samples from Lawrence Livermore lab., four-pass pump relay recover.

 

 

 

   

alignment

 

05/17

Friday

Regen in diagnostic regime , laser maintenance:the test of damage threshold for diffraction grating samples from Lawrence Livermore lab., four-pass pump relay recover.

 

 

 

   

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

05/19

Monday

 

Regen in diagnostic regime, laser maintenance: four-pass pump relay alignment

wavefront correction alignment

 

pump alignment

05/20

Tuesday

 

 

Regen in diagnostic regime, two-pass amplifier pump cavity recover

 

wavefront correction alignment

 

pump alignment

05/21

Wednesday

 

Regen in diagnostic regime, two-pass amplifier pump cavity recover , beam diagnostic system alignment

 

wavefront correction alignment

 

faraday rotator recover

 

05/22

Thursday

Regen in diagnostic regime , laser maintenance, two-pass amplifier pump cavity recover, two-pass pump relay recover, beam diagnostic system alignment

 

wavefront correction alignment

 

   

 

 

05/23

Friday

Regen in diagnostic regime.

 

wavefront correction alignment

 

   

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

05/27

Monday

 

holiday

 

 

 

05/28

Tuesday

 

 

Regen in diagnostic regime,

 

wavefront correction alignment

 

 

05/29

Wednesday

 

Regen in full power, cleaner on, four-pass on, two-pass on

 

 

firing with 10TW

f

 

05/30

Thursday

Regen in diagnostic regime

 

wavefront correction alignment

 

   

 

 

05/31

Friday

Regen in diagnostic regime.

 

wavefront correction alignment

 

   

 

 

JUNE 2003

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

06/03

Monday

 

Regen in diagnostic regime, spider alignment

 

 

 

06/04

Tuesday

 

 

Chiller is out, the reflector of two-pass amplifier recover

 

 

 

 

06/05

Wednesday

 

Chiller is out, the reflector of two-pass amplifier recover

 

 

 

 

 

06/06

Thursday

Chiller is out, the reflector of two-pass amplifier recover

 

   

 

 

05/07

Friday

Regen in diagnostic regime, spider alignment

 

   

 

 

 

DATE  

HERCULES    

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

06/09

Monday

 

Regen in diagnostic regime, spider alignment

 

 

 

06/10

Tuesday

 

 

Regen in diagnostic regime, spider alignment

 

 

 

 

06/11

Wednesday

 

FOCUS meeting

 

 

 

 

 

06/12

Thursday

FOCUS meeting

 

   

 

 

05/13

Friday

Regen in diagnostic regime, spider alignment, two-pass amplifier recover

 

   

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

06/16

Monday

Regen in diagnostic regime, far-field measurements to check the source of laser instability

 

 

 

06/17

Tuesday

 

 

Regen in diagnostic regime, far-field measurements to check the source of laser instability

 

 

 

06/18

Wednesday

 

Regen in full power

 

hole drilling

 

pump

 alignment

 

06/19

Thursday

Regen in full power,

 

hole drilling

 

   

 

 

06/20

Friday

Regen in diagnostic regime for spider alignment

 

   

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

06/23

Monday

 

Oscillator repair, the installation of the crystal for 100TW setup

 

 

pump

 alignment

 

06/24

Tuesday

 

 

Oscillator repair, the installation of the crystal for 100TW setup

 

 

 

06/25

Wednesday

 

Oscillator repair, the installation of the crystal for 100TW setup

 

 

 

06/26

Thursday

Oscillator repair, the installation of the crystal for 100TW setup

 

 

   

 

 

06/27

Friday

Oscillator repair, the installation of the crystal for 100TW setup

 

   

 

 

 

July 2003

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

06/30

Monday

 

Oscillator repair, the installation of the crystal for 100TW setup

 

 

 

07/01

Tuesday

 

 

A diode for oscillator pump has been installed, the new coded crystal for 100TW setup has been installed, the crystal has been pumped with long and short pulses

 

 

 

07/02

Wednesday

 

The diode for oscillator pump alignment, oscillator chiller repair

 

 

 

07/03

Thursday

Oscillator system alignment

 

 

   

Pump alignment in short pulse mode

 

07/04

Friday

holiday

 

   

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

07/07

Monday

 

Pockels call #2 repair

 

 

 

Pump alignment in short pulse mode

07/08

Tuesday

 

 

Pockels call #2 repair and alignment

 

 

 

Pump alignment in short pulse mode

07/09

Wednesday

 

Pockels call #3 alignment

 

 

 

Pump alignment in short pulse mode

07/10

Thursday

Regen in diagnostic regime, regen in full power

 work with solid target(alignment)

 

user  refused

 

Pump alignment in short pulse mode

 

07/11

Friday

Regen in diagnostic regime, regen in full power

 work with solid target(alignment)

 

user  refused

 

 

Pump alignment in short pulse mode

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

07/13

Monday

 

Regen in diagnostic regime

 

work with solid target (alignment)

 

Pump alignment in short pulse mode, pump channel building

07/14

Tuesday

 

 

Regen in diagnostic regime, regen in full power

 

work with solid target (alignment)

 

Pump alignment in short pulse mode, pump channel building

07/15

Wednesday

 

Regen in full power

 

 

work with solid target (alignment)

 

 

Pump alignment in short pulse mode, pump channel building

07/16

Thursday

Regen in diagnostic regime, regen in full power

work with solid target (alignment)

 

 

 

Pump alignment in short pulse mode

 

07/17

Friday

Regen in diagnostic regime, regen in full power,cleaner on , four-pass on

study of preplasma formation

 

 

 

 

Pump alignment in short pulse mode

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC

PETAWATT

07/21

Monday

Crystal of two-pass amplifier has been changed and aligned. Regen in diagnostic regime

 

alignment

Pump alignment in short pulse mode, pump channel building

07/22

Tuesday

 

 

Regen in diagnostic regime, regen in full power

alignment 

 

Pump alignment in short pulse mode, pump channel building

07/23

Wednesday

Regen in diagnostic regime, regen in full power

 

alignment

 

 

Pump alignment in short pulse mode, pump channel building

07/24

Thursday

Regen in diagnostic regime, regen in full power, cleaner on, 4-pass on, 2-pass on

 

 

firing with 30 TW

 

Pump alignment in short pulse mode

 

07/25

Friday

vocation

vocation

 

vocation

 

 

vocation

 

 

July 25 - August 3    VOCATION!!!

August 2003

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

08/05

Tuesday

 

Regen in diagnostic regime, oscillator and regen alignment

     

08/06

Wednesday

 

Regen in diagnostic regime, regen in full power, 4 - pass amplifier alignment to improve a beam profile, water filters installation for chilling systems of PRO and Macho Light lasers

     

08/07

Thursday

 

Regen in diagnostic regime, regen in full power, compressor chamber alignment to make measurements of pulse duration and spectrum in 30 TW regime, water filters installation for chilling systems of Glass laser and Petawatt pump laser

     

08/08

Friday

Regen in diagnostic regime, problems with chilling system of regen Pockels call, maintenance day

     

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

08/11

Monday

 

Regen chiller maintenance, regen alignment      

08/12

Tuesday

 

Regen in diagnostic regime, regen in full power, four-pass amp. on

 

   

08/13

Wednesday

 

Regen in diagnostic regime,

second order autocorrelator alignment, oscillator in Macho Light laser alignment  

 

   

08/14

Thursday

 

Regen in diagnostic regime

Switching-off of an electricity

solid target alignment

   

08/15

Friday

no electricity, Gerard Mourou brought his own power generator to run ion pumps

     

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

08/18

Monday

Regen in diagnostic regime rarified gas target alignment    

08/19

Tuesday

 

Regen in diagnostic regime, regen in full power

rarified gas target alignment

   

08/20

Wednesday

 

Regen in diagnostic regime

regen in full power

rarified gas target alignment

   

08/21

Thursday

 

Regen in diagnostic regime

regen in full power

 

rarified gas target alignment

   

08/22

Friday

Regen in diagnostic regime

regen in full power

firing with energy from 5 up to 15 mJ into rarified gas target    

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

08/25

Monday

Regen in diagnostic regime,

regen in full power 

firing with pulse energy up to 7  mJ into rarified gas target, the goal of the experiment is to observe electrons after noble gas atoms ionization by laser light.    

08/26

Tuesday

 

Regen in diagnostic regime, regen in full power 

firing with pulse energy up to 7  mJ into rarified gas target. No electrons has been detected.

   

08/27

Wednesday

 

Regen in diagnostic regime,

regen in full power 

firing with pulse energy up to 7  mJ into rarified gas target. No electrons has been detected.

   

08/28

Thursday

 

Regen in diagnostic regime, regen in full power, Macho Light laser maintenance      

08/29

Friday

Regen in diagnostic regime, regen in full power

firing with pulse energy up to 7  mJ into rarified gas target. No electrons has been detected.    

 

September 2003

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

09/01

Monday

LABOR DAY

LABOR DAY  

LABOR DAY

 

 

LABOR DAY

 

09/02

Tuesday

 

Regen in diagnostic regime, regen in full power, 4-pass amp. on, 2-pass amp. on. No amplification in 2-pass amp. setup because crystal position is wrong

solid target alignment.

   

09/03

Wednesday

 

Regen in diagnostic regime, 2-pass amp. crystal has been installed proper way

solid target alignment with diode laser

   

09/04

Thursday

 

Regen in diagnostic regime

solid target alignment with diode laser

   

09/05

Friday

Regen in diagnostic regime, regen in full power, 4-pass amp. on. Bad beam profile after 4-pass has been found out. The reason is  - old charging lamps in amplifier and preamplifier of Macho Light laser (4-pass pump)

solid target alignment with diode laser    

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

09/08

Monday

Regen in diagnostic regime, lamps in preamplifier and both  amplifiers of Macho Light laser (4- pass amplifier pump) has been changed 

solid target alignment    

09/09

Tuesday

 

High voltage power supply in Regen pump (PRO laser) is broken, new one has been ordered from Spectra Physics.   Macho Light laser maintenance.

solid target alignment with diode laser

   

09/10

Wednesday

 

High voltage power supply in Regen pump (PRO laser) is broken, new one has been ordered from Spectra Physics.   Macho Light laser maintenance.

solid target alignment with diode laser

 

Yellow channel of petawatt pump alignment

09/11

Thursday

 

 

solid target alignment with diode laser

 

Yellow channel of petawatt pump alignment, the rod in yellow head has been burned

09/12

Friday

 

   

An extra rod cleaning to put instead burned one

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

09/15

Monday

The power supply installation into PRO amplifier , Regen in diagnostic regime, regen in full power, 4-pass on

solid target alignment.

 

 

An extra rod cleaning to put instead burned one

09/16

Tuesday

 

Regen in diagnostic regime

solid target alignment.

  New rod installation into yellow head of petawatt pump

09/17

Wednesday

 

Regen in diagnostic regime

solid target alignment

 

 

09/18

Thursday

 

Regen in full power

drilling holes in the solid target

 

 

09/19

Friday

Regen in full power

drilling holes in the solid target

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

10/20

Monday

Regen in diagnostic regime

solid target alignment.

 

 

 

10/21

Tuesday

 

Regen in full power

drilling holes in the solid target

   

10/22

Wednesday

 

Regen in full power

drilling holes in the solid target

 

 

 

10/23

Thursday

 

Regen in diagnostic regime,

4 - pass amp. alignment

 

 

 

10/24

Friday

Regen in diagnostic regime 4 - pass amp. alignment , water pump in Macho Light power supply was broken

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

10/27

Monday

Regen in diagnostic regime

solid target alignment.

 

 

 

10/28

Tuesday

 

Regen in full power

drilling holes in the solid target

   

10/29

Wednesday

 

Regen in full power, new water pump installation into Macho Light power supply

drilling holes in the solid target

 

 

 

10/30

Thursday

 

Regen in diagnostic regime, Regen in long pulse, 4 - pass amp. on, checking of the beam profile at the exit of 4-pass

drilling holes in the solid target

 

 

10/31

Friday

Regen in full power

drilling holes in the solid target

 

 

 

November 2003

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

11/03

Monday

Regen in full power

drilling holes in the solid target

 

 

 

11/04

Tuesday

 

Regen in full power

drilling holes in the solid target

   

11/05

Wednesday

 

Oscillator problems because vibration

 

 

 

 

11/06

Thursday

 

Oscillator alignment

 

 

 

11/07

Friday

Regen in full power

drilling holes in the solid target

 

 

 

 

DATE

HERCULES

ABOVE THRESHOLD IONIZATION EXPERIMENT

LILAC PETAWATT

11/10

Monday

Regen in full power

drilling holes in the solid target

 

 

 

11/11

Tuesday

Regen in full power

drilling holes in the solid target

   

11/12

Wednesday

Regen in full power 

drilling holes in the solid target

 

 

11/13

Thursday

Regen in full power 

drilling holes in the solid target

 

 

11/14

Friday

Regen in full power

drilling holes in the solid target

 

 

Research Index / Hercules / Schedule of previous weeks / 2003

Copyright © Center for Ultrafast Optical Science, University of Michigan

Maintained by: cuos-webmaster@umich.edu